Follow the seven step strategy to graph the following rational function. A. The x-intercept(s) is/are 0 (Type an integer or a i need to find the last one
Answer #1

y = +9

take x=-2

2(-2) y=-2) +9

2.4 9 4+9

دان

.

take x=-1

| 2 (-1)2 y = (-10 +9

y = 2.1 1+9

y=\frac{2}{10}

{\color{Red} y=\frac{1}{5}}

.

.

take x=5

2 (5) 9+ (ت = 9

2.25 25 +9

y=\frac{50}{34}

{\color{red} y=\frac{25}{17}}

.

.

take x=6

6 = 9اج +6 7 (9)ج

y=\frac{2\cdot 36}{36+9}

y=\frac{72}{45}

|ا

.

make final table

x -2 -1 5 6
y = +9 {\color{Red} \frac{8}{13}} {\color{Red} \frac{1}{5}} {\color{red} \frac{25}{17}} {\color{Red} \frac{8}{5}}